OpenStudy (anonymous):

order of operationr simplify the expression 5/12 - 1/2 X 1/6

OpenStudy (anonymous):

5/12 - 1/12 = 4/12 = 1/3

OpenStudy (anonymous):

but where is the 1/6 in there?

OpenStudy (anonymous):

1/2*1/6 = 1/12

OpenStudy (anonymous):

i dont see where you get 1/12 from

OpenStudy (anonymous):

1/2 X 1/6 = 1/12 Multiplication goes first in order of operations

OpenStudy (anonymous):

ok so what is the final answer for 5/12- 1/2 X 1/6

OpenStudy (anonymous):

1/3